OpenStudy (anonymous):

Find the solutions x^2+4x+4=9

OpenStudy (anonymous):

x^2 + 4x -5 = 0 (x+5)(x-1) = 0 x = 1 and x = -5

OpenStudy (anonymous):

x^2 + 4x + 4 = 9 4 + 4x + x2 = 9 4 + 4x + x2 = 9 -9 =-9 --------------- -5 + 4x + x2 = 9 + -9 -5 + 4x + x2 = 0 (-5 + -1x)(1 + -1x) = 0 ==================================== -5 + -1x = 0 -5 + 5 + -1x = 0 + 5 0 + -1x = 0 + 5 -1x = 0 + 5 -1x = 5 Divide by -1 x = -5 1 + -1x = 0 1 + -1 + -1x = 0 + -1 0 + -1x = 0 + -1 -1x = 0 + -1 -1x = -1 Divide each side by -1 x = 1 ========================== Solution x = {-5, 1}